# On local Galois deformation rings: generalised reductive groups

### v1

@@ 1.7. Notation
Let $L$ be another finite extension of $\mathbb{Q}_{p}$, such that $\operatorname{Hom}_{\mathbb{Q}_{p}\text{-}\mathrm{alg}}(F,L)$ has cardinality $[F:\mathbb{Q}_{p}]$. Let $\mathcal{O}$ be the ring of integers in $L$, $\varpi$ a uniformiser, and $k$ the residue field. All our schemes are defined over $\mathcal{O}$ (unless stated otherwise). We omit $\operatorname{Spec}\mathcal{O}$ from the notation, when we take fibre products over it, so that $X\times Y:=X\times_{\operatorname{Spec}\mathcal{O}}Y$.
@@
be its linearisation.
We denote by $A\text{-}\mathrm{alg}$ the category of commutative $A$-algebras. For a commutative ring $R$ we let $P_{1}R=\{\mathfrak{p}\in\operatorname{Spec}R:\dim R/\mathfrak{p}=1\}$.
@@ 1.8. Acknowledgements
The authors would like to thank Jarod Alper and Sean Cotner for the correspondence regarding their work; Tasho Kaletha for a very useful discussion regarding Section 11; Vincent Pilloni for sharing his notes of the talk he gave on [8]; Sophie Morel for making a preliminary version of [25] available to them; Toby Gee, James Newton and Sean Cotner for their comments. The authors would like to especially thank Brian Conrad for pointing out an error in an earlier version of the paper and for the subsequent correspondence.
@@ 2. $\mathcal{O}$-group schemes
The category of $\mathcal{O}$-group schemes embeds fully faithfully in the category of fppf group sheaves on $(\mathcal{O}\text{-}\mathrm{alg})^{\mathrm{op}}$. Whenever an exact sequence is considered it is understood as an exact sequence of fppf group sheaves. If $G$ is a flat $\mathcal{O}$-group scheme of finite presentation and $N\leq G$ is a flat closed normal subgroup, then the quotient group sheaf $G/N$ is representable by a scheme [4, Théorème 4.C]. If $G$ and $N$ are generalised reductive then the quotient is affine by [3, Theorem 9.4.1].
@@
This is a special case of [1, Exposé VIII, Théorème 6.4] as explained in [1, Exposé VIII, Exemples 6.5 (e)]. ∎
@@
Since $G_{1}$ is flat and of finite presentation, the map of $\mathcal{O}$-schemes
$G_{1}\times G_{2}\to G_{2}\times G_{2},~{}(g,h)\mapsto(gh,h)$
is an fppf equivalence relation. The projection map onto the second factor is clearly faithfully flat and of finite presentation. The first projection map is faithfully flat and of finite presentation, since multiplication is isomorphic to projection onto the second factor, as can be seen by considering the automorphism $(g,h)\mapsto(g,g^{-1}h)$ of $G_{1}\times G_{2}$.
By definition of exactness $G_{3}$ is the quotient sheaf of $G_{2}$ by this equivalence relation. Since $G_{3}$ is already representable by a scheme, we know that the equivalence relation is effective by [30, Exposé IV, Corollaire 6.3.3]. By definition of effectivity [30, Exposé IV, Déf. 3.4.3] this means, that the map $G_{2}\to G_{3}$ is faithfully flat and of finite presentation. ∎
@@
It follows from [20, Proposition 3.1.3], that for a generalised reductive $\mathcal{O}$-group scheme $G$ the quotient $G/G^{0}$ is étale. Sean Cotner has shown in [22] using the results of Alper that a smooth affine $S$-group scheme $G$ is generalised reductive if and only if it is geometrically reductive in the sense of [3, Definition 9.1.1]. In particular, this holds over $S=\operatorname{Spec}\mathcal{O}$.
@@
The first assertion follows from [54, Tag 047N, Tag 047P]. Since $\kappa$ is perfect, if the unipotent radical of $G$ is trivial then also the unipotent radical of $G_{K}$ is trivial, for any field extension $K$ of $\kappa$ by [21, Proposition 1.1.9]. Thus $G^{0}$ is reductive and $G/G^{0}$ is finite by [20, Proposition 3.1.3]. ∎
@@
By definition $G/G^{0}$ is finite étale. Using [54, Tag 0BND] and the fact that the étale fundamental group of $\mathcal{O}$ is the Galois group of the maximal unramified extension of $L$ we find a finite unramified extension $L^{\prime}$, such that $(G/G^{0})_{\mathcal{O}^{\prime}}$ is constant. Since the extension $\mathcal{O}\rightarrow\mathcal{O}^{\prime}$ is flat the sequence of group schemes $0\rightarrow G^{0}\rightarrow G\rightarrow G/G^{0}\rightarrow 0$ remains exact after base change to $\mathcal{O}^{\prime}$.
@@
By [20, Theorem 5.3.1] $G^{\prime}$ is a semisimple $\mathcal{O}$-group scheme and in particular smooth. Since $G^{\prime}$ is the fppf sheafification of the commutator subgroup functor $A\mapsto[G^{0}(A),G^{0}(A)]$, it is preserved under any automorphism of $G^{0}$. So $G^{\prime}$ is a closed normal subgroup scheme of $G$. By 2.3 the map $\pi$ is faithfully flat and of finite presentation. So by [54, Tag 01V8] we need to see, that all fibres of $\pi$ are smooth. Let $\kappa$ be any field and let $x:\operatorname{Spec}(\kappa)\to G/G^{\prime}$ be a $\kappa$-point; this also defines a map $\mathcal{O}\to\kappa$. So we get a natural map $x^{\prime}:\operatorname{Spec}(\kappa)\to(G/G^{\prime})_{\kappa}$, where $(G/G^{\prime})_{\kappa}=G/G^{\prime}\times\operatorname{Spec}(\kappa)$. The fibre of $\pi$ above $x$ coincides with the fibre of $\pi_{\kappa}:G_{\kappa}\to(G/G^{\prime})_{\kappa}$ above $x^{\prime}$. By translation in $(G/G^{\prime})_{k}$ this fibre is isomorphic to $G_{\kappa}^{\prime}$, which is smooth. ∎
@@ 2.1. R-parabolics
We recall some facts about parabolic subgroups of a generalised reductive $\mathcal{O}$-group scheme $G$. The main issue is that we do not assume that $G$ is connected. The letter R stands for Richardson. Let $A$ be an $\mathcal{O}$-algebra and let $\lambda:\mathbb{G}_{m,A}\rightarrow G_{A}$ be a cocharacter. We then define $P_{\lambda}$ (resp. $U_{\lambda}$) to be the closed $A$-group scheme of $G_{A}$ representing the functor of points $g$ of $G$ such that the limit $\lim_{t\rightarrow 0}\lambda(t)g\lambda(t)^{-1}$ exists (resp. exists and is equal to the identity), see [21, Lemmas 2.1.4, 2.1.5]. Let $L_{\lambda}=P_{\lambda}\cap P_{-\lambda}$, where $-\lambda$ is the cocharacter defined by $-\lambda(t):=\lambda(1/t)=\lambda(t)^{-1}$. We call subgroup schemes of $G$ the form $P_{\lambda}$ R-parabolic subgroup schemes and $L_{\lambda}$ R-Levi subgroup schemes of $P_{\lambda}$. We will only consider R-parabolic and R-Levi subgroup schemes over a connected base.
@@
Let $\lambda:\mathbb{G}_{m}\rightarrow G$ be a cocharacter such that $P=P_{\lambda,G}$ and $L=L_{\lambda,G}$. Since $\mathbb{G}_{m}$ is connected $\lambda$ factors through $\lambda:\mathbb{G}_{m}\rightarrow G^{0}$. It follows from [21, Lemma 2.1.5] that $P\cap G^{0}=P_{\lambda,G^{0}}$ and $L\cap G^{0}=L_{\lambda,G^{0}}$. Let $\bar{x}$ be a geometric point of $\operatorname{Spec}\mathcal{O}$. Then $(P_{\lambda,G^{0}})_{\bar{x}}=P_{\lambda_{\bar{x}},G^{0}_{\bar{x}}}$ and $(L_{\lambda,G^{0}})_{\bar{x}}=L_{\lambda_{\bar{x}},G^{0}_{\bar{x}}}$. It follows from [21, Proposition 2.2.9] that $P_{\lambda_{\bar{x}},G^{0}_{\bar{x}}}$ is a parabolic subgroup of $G^{0}_{\bar{x}}$ with Levi $L_{\lambda_{\bar{x}},G^{0}_{\bar{x}}}$. Moreover, both groups are connected and $L_{\lambda_{\bar{x}},G^{0}_{\bar{x}}}$ is reductive by classical theory of reductive groups over algebraically closed fields. Thus $P\cap G^{0}$ is the neutral component of $P$, $L\cap G^{0}$ is the neutral component of $L$. Moreover, $P^{0}$ is a parabolic subgroup of $G^{0}$ in the sense of [20, Definition 5.2.1] and it follows from Lemma 2.12 (3) that $L^{0}$ is a Levi of $P^{0}$ in the sense of [20, Definition 5.4.2]. Further, $L^{0}$ is reductive in the sense of [20, Definition 3.1.1]. The map $L\rightarrow G/G^{0}$ identifies $L/L^{0}$ with a closed subgroup of $G/G^{0}$. Hence $L/L^{0}$ is finite over $\mathcal{O}$. We deduce that $L$ is generalised reductive and $L/L^{0}$ is finite étale over $\mathcal{O}$ by Remark 2.6.
@@
We will first establish that $P\mapsto P_{\bar{s}}$ induces the bijection between sets in parts (1) and (3). If $\lambda:\mathbb{G}_{m}\rightarrow G$ is a cocharacter defined over $\mathcal{O}$ then $(P_{\lambda})_{\bar{s}}=P_{\lambda_{\bar{s}}}$ by [21, Lemma 2.1.4]. Moreover, if $P$ contains $T$ then $P_{\bar{s}}$ will contain $T_{\bar{s}}$. Hence, the map is well defined.
If $P$ and $Q$ are R-parabolics of $G$ such that $P_{\bar{s}}=Q_{\bar{s}}$ then $P^{0}_{\bar{s}}=Q^{0}_{\bar{s}}$. It follows from [20, Corollary 5.2.7 (2)] that there exists a Zariski open subset $U$ of $\operatorname{Spec}\mathcal{O}$ containing $s$ such that $P^{0}_{U}=Q^{0}_{U}$. Since $\mathcal{O}$ is a DVR we have $U=\operatorname{Spec}\mathcal{O}$ and hence $P^{0}=Q^{0}$. It follows from Lemma 2.13 that $P\cap G^{0}=P^{0}=Q^{0}=Q\cap G^{0}$. Thus it is enough to show that the images of $P$ and $Q$ in $G/G^{0}$ coincide. Since $G/G^{0}$ is a constant group scheme, we may check this after base change to $\bar{s}$, which holds as $P_{\bar{s}}=Q_{\bar{s}}$.
Let $\bar{\lambda}:\mathbb{G}_{m}\rightarrow G_{\bar{s}}$ be a cocharacter. Its image is contained in a maximal split torus $T^{\prime}$ of $G_{\bar{s}}$. Moreover, $T^{\prime}$ is contained in $P_{\bar{\lambda}}$. If $P_{\bar{\lambda}}$ contains $T_{\bar{s}}$ then there is $g\in P_{\bar{\lambda}}(\kappa(\bar{s}))$ such that $T_{\bar{s}}=gT^{\prime}g^{-1}$ by [11, Proposition 11.19]. We thus may assume that the image of $\bar{\lambda}$ is contained in $T_{\bar{s}}$. Since both $\mathbb{G}_{m}$ and $T$ are split tori defined over $\mathcal{O}$ there is a cocharacter $\lambda:\mathbb{G}_{m}\rightarrow T$, such that $\lambda_{\bar{s}}=\bar{\lambda}$. Then $P_{\lambda}$ contains $T$ and satisfies $(P_{\lambda})_{\bar{s}}=P_{\bar{\lambda}}$.
@@
If $L$ and $M$ are R-Levi subgroups of $P$ containing $T$ then $M_{\bar{x}}=L_{\bar{x}}$ for every geometric point $\bar{x}$ of $\operatorname{Spec}\mathcal{O}$ by [5, Corollary 6.5]. This implies that $L=M$. ∎
@@
Let $\lambda:\mathbb{G}_{m,\bar{x}}\rightarrow G_{\bar{x}}$ be a cocharacter such that $Q=P_{\lambda}$ and $M=L_{\lambda}$. There is a maximal torus $T^{\prime}$ of $G_{\bar{x}}$, such that the image of $\lambda$ is contained in $T^{\prime}$. Since tori are connected, both $T^{\prime}$ and $T_{\bar{x}}$ are contained in $G^{0}_{\bar{x}}$. Since $\kappa(\bar{x})$ is algebraically closed any two maximal tori in $G^{0}_{\bar{x}}$ are conjugate, and hence after conjugating $Q$ we may assume that $\lambda:\mathbb{G}_{m,\bar{x}}\rightarrow T_{\bar{x}}$. Since $T$ is split over $\mathcal{O}$, there is a cocharacter $\mu:\mathbb{G}_{m}\rightarrow T$ defined over $\mathcal{O}$ such that $\lambda=\mu_{\bar{x}}$. We have $P_{\mu}\times_{\operatorname{Spec}\mathcal{O}}\bar{x}=P_{\mu_{\bar{x}}}$ by [21, Lemma 2.1.4] and $L_{\mu}\times_{\operatorname{Spec}\mathcal{O}}\bar{x}=L_{\mu_{\bar{x}}}$ by [21, Lemma 2.1.5]. ∎
2.2. $G$-semisimplification
@@
If $G$ is connected then this is proved in [51, Proposition 3.3 (b)]. The same argument is carried out in our more general setting in [44, Proposition 2.15]. ∎
@@
In [51, Section 3.2.1] Serre introduced the notion of $G$-complete reducibility in the connected case. We extend this definition to generalised reductive group schemes verbatim.
@@
Let $P$ be a minimal R-parabolic, such that $\rho(\Gamma)\subseteq P(\kappa)$. If $\rho$ is $G$-completely reducible, there exists some R-Levi $L$ of $P$, such that $\rho(\Gamma)\subseteq L(\kappa)$. In particular $\rho=c_{P,L}\circ\rho$. We can apply 2.20 to conclude, $\rho^{\mathrm{ss}}$ is the $G^{0}(\kappa)$-conjugacy class of $\rho$. Conversely, suppose that $\rho\in\rho^{\mathrm{ss}}$, let $L$ be any R-Levi of $P$. The assumption implies that $\rho$ is conjugate to $c_{P,L}\circ\rho$ by an element of $g\in G^{0}(\kappa)$. Minimality of $P$ and [5, Corollary 6.4] imply that the image of $c_{P,L}\circ\rho$ is $L$-irreducible, which means that it is not contained in any proper R-parabolic of $L$. Hence, $\rho(\Gamma)$ is $gLg^{-1}$-irreducible. It follows from the equivalence of parts (iv) and (v) of [5, Corollary 3.5], which as explained in [5, Section 6.3] also holds for non-connected groups, that $\rho(\Gamma)$ is $G$-completely reducible. ∎
@@
In general $(\tau\circ\rho)^{\mathrm{ss}}$ is not isomorphic to $\tau\circ\rho^{\mathrm{ss}}$. But it follows from 2.24, that $D_{\tau\circ\rho^{\mathrm{ss}}}=D_{\tau\circ\rho}$ and in particular from [16, Theorem 2.12] that $(\tau\circ\rho^{\mathrm{ss}})^{\mathrm{ss}}$ is isomorphic to $(\tau\circ\rho)^{\mathrm{ss}}$.
@@
Let $\lambda,\mu:\mathbb{G}_{m}\rightarrow G$ be cocharacters such that $L=L_{\lambda}$ and $M=L_{\mu}$. Since $H$ is contained in $L$ and in $M$, Lemma 2.12 (1) implies that the both $\lambda$ and $\mu$ factor through the inclusion $Z_{G}(H)\subseteq G$. Let $S$ (resp. $T$) be a maximal torus of $Z_{G}(H)$ containing the image of $\lambda$ (resp. $\mu$). It follows from the proof of [11, Proposition 8.18] that $Z_{G}(S)$ is an R-Levi of $G$ containing $H$ and contained in $L$. Since $H$ is $L$-irreducible by assumption we get that $L=Z_{G}(S)$. Similarly, we obtain that $M=Z_{G}(T)$. There is $g\in Z_{G}(H)^{0}(\kappa)$ such that $S=gTg^{-1}$ by [11, Proposition 11.19]. Thus $L=Z_{G}(gTg^{-1})=gZ_{G}(T)g^{-1}=gMg^{-1}$. ∎

### v2

@@ 1.7. Notation
Let $L$ be another finite extension of $\mathbb{Q}_{p}$, such that $\operatorname{Hom}_{\mathbb{Q}_{p}\text{-}\mathrm{alg}}(F,L)$ has cardinality $F:\mathbb{Q}_{p}]$. Let $\mathcal{O}$ be the ring of integers in $L$, $\varpi$ a uniformiser, and $k$ the residue field. All our schemes are defined over $\mathcal{O}$ (unless stated otherwise). We omit $\operatorname{Spec}\mathcal{O}$ from the notation, when we take fibre products over it, so that $X\times Y:=X\times_{\operatorname{Spec}\mathcal{O}}Y$.
@@
be its linearisation, where $M_{d}(A)$ is the set of $d\times d$ matrices with entries in $A$.
We denote by $A\text{-}\mathrm{alg}$ the category of commutative $A$-algebras.
@@ 1.8. Acknowledgements
The authors would like to thank Jarod Alper and Sean Cotner for the correspondence regarding their work; Tasho Kaletha for a very useful discussion regarding Section 11; Vincent Pilloni for sharing his notes of the talk he gave on [6]; Sophie Morel for making a preliminary version of [25] available to them; Toby Gee, James Newton and Sean Cotner for their comments. The authors would like to especially thank Brian Conrad for pointing out an error in an earlier version of the paper and for the subsequent correspondence. We also heartily thank the referee for their careful reading of the paper.
@@ 2. $\mathcal{O}$-group schemes
The category of $\mathcal{O}$-group schemes embeds fully faithfully in the category of fppf group sheaves on $(\mathcal{O}\text{-}\mathrm{alg})^{\mathrm{op}}$. Whenever an exact sequence is considered it is understood as an exact sequence of fppf group sheaves. If $G$ is a flat $\mathcal{O}$-group scheme of finite presentation and $N\leq G$ is a flat closed normal subgroup, then the quotient group sheaf $G/N$ is representable by a scheme [2, Théorème 4.C]. If $G$ and $N$ are generalised reductive then the quotient is affine by [1, Theorem 9.4.1].
@@
This is a special case of [22, Exposé VIII, Théorème 6.4] as explained in [22, Exposé VIII, Exemples 6.5 (e)]. ∎
@@
This is [30, Exposé VIB, Proposition 9.2 (xi)]. ∎
@@
It follows from [20, Proposition 3.1.3], that for a generalised reductive $\mathcal{O}$-group scheme $G$ the quotient $G/G^{0}$ is étale. Sean Cotner has shown in [21] using the results of Alper that a smooth affine $S$-group scheme $G$ is generalised reductive if and only if it is geometrically reductive in the sense of [1, Definition 9.1.1]. In particular, this holds over $S=\operatorname{Spec}\mathcal{O}$.
@@
The first assertion follows from [59, Tag 047N, Tag 047P]. Since $\kappa$ is perfect, if the unipotent radical of $G$ is trivial then also the unipotent radical of $G_{K}$ is trivial, for any field extension $K$ of $\kappa$ by [16, Proposition 1.1.9]. Thus $G^{0}$ is reductive and $G/G^{0}$ is finite by [20, Proposition 3.1.3]. ∎
@@
By definition $G/G^{0}$ is finite étale. Using [59, Tag 0BND] and the fact that the étale fundamental group of $\mathcal{O}$ is the Galois group of the maximal unramified extension of $L$ we find a finite unramified extension $L^{\prime}$, such that $(G/G^{0})_{\mathcal{O}^{\prime}}$ is constant. Since the extension $\mathcal{O}\rightarrow\mathcal{O}^{\prime}$ is flat the sequence of group schemes $0\rightarrow G^{0}\rightarrow G\rightarrow G/G^{0}\rightarrow 0$ remains exact after base change to $\mathcal{O}^{\prime}$.
@@
By [20, Theorem 5.3.1] $G^{\prime}$ is a semisimple $\mathcal{O}$-group scheme and in particular smooth. Since $G^{\prime}$ is the fppf sheafification of the commutator subgroup functor $A\mapsto[G^{0}(A),G^{0}(A)]$, it is preserved under any automorphism of $G^{0}$. So $G^{\prime}$ is a closed normal subgroup scheme of $G$. By 2.3 the map $\pi$ is faithfully flat and of finite presentation. By [30, Exposé VIB, Proposition 9.2 (vii)] smoothness of $G^{\prime}$ implies smoothness of $\pi$. ∎
@@ 2.1. R-parabolics
We recall some facts about parabolic subgroups of a generalised reductive $\mathcal{O}$-group scheme $G$. The main issue is that we do not assume that $G$ is connected. The letter R stands for Richardson. Let $A$ be an $\mathcal{O}$-algebra and let $\lambda:\mathbb{G}_{m,A}\rightarrow G_{A}$ be a cocharacter. We then define $P_{\lambda}$ (resp. $U_{\lambda}$) to be the closed $A$-group scheme of $G_{A}$ representing the functor of points $g$ of $G$ such that the limit $\lim_{t\rightarrow 0}\lambda(t)g\lambda(t)^{-1}$ exists (resp. exists and is equal to the identity), see [16, Lemmas 2.1.4, 2.1.5]. Let $L_{\lambda}=P_{\lambda}\cap P_{-\lambda}$, where $\lambda$ is the cocharacter defined by $\lambda(t):=\lambda(1/t)=\lambda(t)^{-1}$. We call subgroup schemes of $G$ the form $P_{\lambda}$ R-parabolic subgroup schemes and $L_{\lambda}$ R-Levi subgroup schemes of $P_{\lambda}$. We will only consider R-parabolic and R-Levi subgroup schemes over a connected base.
@@
Let $\lambda:\mathbb{G}_{m}\rightarrow G$ be a cocharacter such that $P=P_{\lambda,G}$ and $L=L_{\lambda,G}$. Since $\mathbb{G}_{m}$ is connected $\lambda$ factors through $\lambda:\mathbb{G}_{m}\rightarrow G^{0}$. It follows from [16, Lemma 2.1.5] that $P\cap G^{0}=P_{\lambda,G^{0}}$ and $L\cap G^{0}=L_{\lambda,G^{0}}$. Let $\bar{x}$ be a geometric point of $\operatorname{Spec}\mathcal{O}$. Then $(P_{\lambda,G^{0}})_{\bar{x}}=P_{\lambda_{\bar{x}},G^{0}_{\bar{x}}}$ and $(L_{\lambda,G^{0}})_{\bar{x}}=L_{\lambda_{\bar{x}},G^{0}_{\bar{x}}}$. It follows from [16, Proposition 2.2.9] that $P_{\lambda_{\bar{x}},G^{0}_{\bar{x}}}$ is a parabolic subgroup of $G^{0}_{\bar{x}}$ with Levi $L_{\lambda_{\bar{x}},G^{0}_{\bar{x}}}$. Moreover, both groups are connected and $L_{\lambda_{\bar{x}},G^{0}_{\bar{x}}}$ is reductive by the classical theory of reductive groups over algebraically closed fields. Thus $P\cap G^{0}$ is the neutral component of $P$ and $L\cap G^{0}$ is the neutral component of $L$. Moreover, $P^{0}$ is a parabolic subgroup of $G^{0}$ in the sense of [20, Definition 5.2.1] and it follows from Lemma 2.12 (3) that $L^{0}$ is a Levi of $P^{0}$ in the sense of [20, Definition 5.4.2]. Further, $L^{0}$ is reductive in the sense of [20, Definition 3.1.1]. The map $L\rightarrow G/G^{0}$ identifies $L/L^{0}$ with a closed subgroup of $G/G^{0}$. Hence $L/L^{0}$ is finite over $\mathcal{O}$. We deduce that $L$ is generalised reductive and $L/L^{0}$ is finite étale over $\mathcal{O}$ by Remark 2.6.
@@
We will first establish that $P\mapsto P_{\bar{s}}$ induces the bijection between sets in parts (1) and (3). If $\lambda:\mathbb{G}_{m}\rightarrow G$ is a cocharacter defined over $\mathcal{O}$ then $(P_{\lambda})_{\bar{s}}=P_{\lambda_{\bar{s}}}$ by [16, Lemma 2.1.4]. Moreover, if $P$ contains $T$ then $P_{\bar{s}}$ will contain $T_{\bar{s}}$. Hence, the map is well defined.
If $P$ and $Q$ are R-parabolics of $G$ such that $P_{\bar{s}}=Q_{\bar{s}}$ then $P^{0}_{\bar{s}}=Q^{0}_{\bar{s}}$. It follows from [20, Corollary 5.2.7 (2)] that there exists a Zariski open subset $U$ of $\operatorname{Spec}\mathcal{O}$ containing $s$ such that $P^{0}_{U}=Q^{0}_{U}$. Since $\mathcal{O}$ is a DVR we have $U=\operatorname{Spec}\mathcal{O}$ and hence $P^{0}=Q^{0}$. It follows from Lemma 2.13 that $P\cap G^{0}=P^{0}=Q^{0}=Q\cap G^{0}$. Thus both $P$ and $Q$ are contained in the $G$-normaliser $N_{G}(P^{0})$ of $P^{0}$ and it is enough to show that the images of $P$ and $Q$ in $N_{G}(P^{0})/P^{0}$ coincide. Since $N_{G}(P^{0})\cap G^{0}=P^{0}$ it is enough to show that the images of $P$ and $Q$ in $G/G^{0}$ coincide. Since $G/G^{0}$ is a constant group scheme, we may check this after base change to $\bar{s}$, which holds as $P_{\bar{s}}=Q_{\bar{s}}$.
Let $\bar{\lambda}:\mathbb{G}_{m}\rightarrow G_{\bar{s}}$ be a cocharacter. Its image is contained in a maximal split torus $T^{\prime}$ of $G_{\bar{s}}$. Moreover, $T^{\prime}$ is contained in $P_{\bar{\lambda}}$. If $P_{\bar{\lambda}}$ contains $T_{\bar{s}}$ then there is $g\in P_{\bar{\lambda}}(\kappa(\bar{s}))$ such that $T_{\bar{s}}=gT^{\prime}g^{-1}$ by [10, Proposition 11.19]. We thus may assume that the image of $\bar{\lambda}$ is contained in $T_{\bar{s}}$. Since both $\mathbb{G}_{m}$ and $T$ are split tori defined over $\mathcal{O}$ there is a cocharacter $\lambda:\mathbb{G}_{m}\rightarrow T$, such that $\lambda_{\bar{s}}=\bar{\lambda}$. Then $P_{\lambda}$ contains $T$ and satisfies $(P_{\lambda})_{\bar{s}}=P_{\bar{\lambda}}$.
@@
If $L$ and $M$ are R-Levi subgroups of $P$ containing $T$ then $M_{\bar{x}}=L_{\bar{x}}$ for every geometric point $\bar{x}$ of $\operatorname{Spec}\mathcal{O}$ by [3, Corollary 6.5]. This implies that $L=M$. ∎
@@
Let $\lambda:\mathbb{G}_{m,\bar{x}}\rightarrow G_{\bar{x}}$ be a cocharacter such that $Q=P_{\lambda}$ and $M=L_{\lambda}$. There is a maximal torus $T^{\prime}$ of $G_{\bar{x}}$, such that the image of $\lambda$ is contained in $T^{\prime}$. Since tori are connected, both $T^{\prime}$ and $T_{\bar{x}}$ are contained in $G^{0}_{\bar{x}}$. Since $\kappa(\bar{x})$ is algebraically closed any two maximal tori in $G^{0}_{\bar{x}}$ are conjugate, and hence after conjugating $Q$ we may assume that $\lambda:\mathbb{G}_{m,\bar{x}}\rightarrow T_{\bar{x}}$. Since $T$ is split over $\mathcal{O}$, there is a cocharacter $\mu:\mathbb{G}_{m}\rightarrow T$ defined over $\mathcal{O}$ such that $\lambda=\mu_{\bar{x}}$. We have $P_{\mu}\times_{\operatorname{Spec}\mathcal{O}}\bar{x}=P_{\mu_{\bar{x}}}$ by [16, Lemma 2.1.4] and $L_{\mu}\times_{\operatorname{Spec}\mathcal{O}}\bar{x}=L_{\mu_{\bar{x}}}$ by [16, Lemma 2.1.5]. ∎
$2.$ $G$-semisimplification
@@
If $G$ is connected then this is proved in [56, Proposition 3.3 (b)]. The same argument is carried out in our more general setting in [50, Proposition 2.15]. ∎
@@
In [56, Section 3.2.1] Serre introduced the notion of $G$-complete reducibility in the connected case. This definition extends to generalised reductive group schemes verbatim.
@@
Let $P$ be a minimal R-parabolic, such that $\rho(\Gamma)\subseteq P(\kappa)$. If $\rho$ is $G$-completely reducible, there exists some R-Levi $L$ of $P$, such that $\rho(\Gamma)\subseteq L(\kappa)$. In particular $\rho=c_{P,L}\circ\rho$. We can apply 2.20 to conclude, $\rho^{\mathrm{ss}}$ is in the $G^{0}(\kappa)$-conjugacy class of $\rho$. Conversely, suppose that $\rho\in\rho^{\mathrm{ss}}$, let $L$ be any R-Levi of $P$. The assumption implies that $\rho$ is conjugate to $c_{P,L}\circ\rho$ by an element of $g\in G^{0}(\kappa)$. Minimality of $P$ and [3, Corollary 6.4] imply that the image of $c_{P,L}\circ\rho$ is $L$-irreducible, which means that it is not contained in any proper R-parabolic of $L$. Hence, $\rho(\Gamma)$ is $gLg^{-1}$-irreducible. It follows from the equivalence of parts (iv) and (v) of [3, Corollary 3.5], which as explained in [3, Section 6.3] also holds for non-connected groups, that $\rho(\Gamma)$ is $G$-completely reducible. ∎
@@
In general $(\tau\circ\rho)^{\mathrm{ss}}$ is not isomorphic to $\tau\circ\rho^{\mathrm{ss}}$. But it follows from 2.24, that $D_{\tau\circ\rho^{\mathrm{ss}}}=D_{\tau\circ\rho}$ and in particular from [15, Theorem 2.12] that $(\tau\circ\rho^{\mathrm{ss}})^{\mathrm{ss}}$ is isomorphic to $(\tau\circ\rho)^{\mathrm{ss}}$.
@@
Let $\lambda,\mu:\mathbb{G}_{m}\rightarrow G$ be cocharacters such that $L=L_{\lambda}$ and $M=L_{\mu}$. Since $H$ is contained in $L$ and in $M$, Lemma 2.12 (1) implies that the both $\lambda$ and $\mu$ factor through the inclusion $Z_{G}(H)\subseteq G$. Let $S$ (resp. $T$) be a maximal torus of $Z_{G}(H)$ containing the image of $\lambda$ (resp. $\mu$). It follows from the proof of [10, Proposition 8.18] that $Z_{G}(S)$ is an R-Levi of $G$ containing $H$ and contained in $L$. Since $H$ is $L$-irreducible by assumption we get that $L=Z_{G}(S)$. Similarly, we obtain that $M=Z_{G}(T)$. There is $g\in Z_{G}(H)^{0}(\kappa)$ such that $S=gTg^{-1}$ by [10, Proposition 11.19]. Thus $L=Z_{G}(gTg^{-1})=gZ_{G}(T)g^{-1}=gMg^{-1}$. ∎